OpenStudy (anonymous):

when using the substitution method, which expression for y in the second equation would be substituted in the first equation? Equation 1: 2x-2y = -12 Equation 2: 4x - y = 3

jhonyy9 (jhonyy9):

2x -2y = -12 4x -y =3 ---- y = ?

jhonyy9 (jhonyy9):

wann you cooperate ?

jhonyy9 (jhonyy9):

ok try adding to both sides y and check what you will get ?

jhonyy9 (jhonyy9):

are you here ?

OpenStudy (anonymous):

sorry! yes i am! wait whats substitution method?

jhonyy9 (jhonyy9):

ok so try making what i have wrote to you above

jhonyy9 (jhonyy9):

substitution method -- what mean ? you dont have learned again ?

OpenStudy (anonymous):

this is for online, and the lessons are that descriptive

OpenStudy (anonymous):

aren't*

jhonyy9 (jhonyy9):

ok so 2x -2y = -12 4x -y =3 ---- y = ? add to the both sides y than what will get ?

jhonyy9 (jhonyy9):

on second equation

OpenStudy (anonymous):

4x = 3 + y ?

jhonyy9 (jhonyy9):

yes and now add to both sides (-3)

OpenStudy (anonymous):

4x - 3 = y

jhonyy9 (jhonyy9):

so with this you have got the y= 4x-3 from second equation than now using the substitution method you need just subtituting in place of y inside first equation this y= 4x-3 ok ?

OpenStudy (anonymous):

so 2x - 2 (4x-3) = -12 ?

jhonyy9 (jhonyy9):

yes and continue

OpenStudy (anonymous):

what do i do when i get that answer? plug in x for the second equation ?

jhonyy9 (jhonyy9):

continue please and i will help you understanding all

OpenStudy (anonymous):

i got x = 3 .

jhonyy9 (jhonyy9):

ok and now subtitute this value of x = 3 in place of x inside y = 4x-3

OpenStudy (anonymous):

y = 9

jhonyy9 (jhonyy9):

ok

jhonyy9 (jhonyy9):

do you understand it now ?

OpenStudy (anonymous):

wait what would my answer be then??

jhonyy9 (jhonyy9):

why ? if you check newly you get the right answer what is the y= ? from second equation

OpenStudy (anonymous):

9

jhonyy9 (jhonyy9):

yes but the question was what will be the expression of y from second equation what you need subtituting in place of y in first equation using the substitution method

OpenStudy (anonymous):

is my answer -4x + 3 ?

jhonyy9 (jhonyy9):

why -4x+3 ?

jhonyy9 (jhonyy9):

you need just checking above what you have substituted in place of y inside first equation

OpenStudy (anonymous):

what , this doesn't make sense . and these are my answer choices -4x - 3 -4x + 3 4x - 3 4x + 3

jhonyy9 (jhonyy9):

so above you have got this 4x - 3 = y

jhonyy9 (jhonyy9):

are you here ?

jhonyy9 (jhonyy9):

ok ?